# A Bayesian joint model of multiple longitudinal and categorical outcomes with application to multiple myeloma using permutation-based variable importance

### v2

@@ 1 Introduction
Many authors have extended the standard joint model. For example, some studies modeled multiple longitudinal outcomes and/or competing risks (Hickey et al., 2016; Rué et al., 2017; Rustand et al., 2024), while others incorporated the dependence of spatial random effects into the joint modeling (Martins et al., 2016, 2017; Rappl et al., 2023). There were also proposals that considered categorical longitudinal and survival outcomes (Choi et al., 2015; Chi et al., 2024), two longitudinal submodels, one continuous and one ordinal (Delporte et al., 2024), and joint models of longitudinal continuous and binary data (Wang et al., 2000; Horrocks and van Den Heuvel, 2009; Lu et al., 2016; Zhou et al., 2023). For the latter, to the best of our knowledge, there are currently no extensions available that accommodate more than two categories. This extension introduces at least two additional complexities: (i) more parameters to be estimated, which requires more data, and (ii) imbalanced categories, leading to poor predictive performance (Kolo, 2010).
@@
The adequacy of the bi-exponential submodels is verified through an analysis of individual weighted residuals (Desmée et al., 2017), while performance of the categorical submodel is evaluated using metrics for multiclass classification (Grandini et al., 2020). Moreover, as a complement to posterior inference, we present a model-agnostic variable importance approach with the aim of improving the transparency, trustability, and interpretability of the results (Casalicchio et al., 2019).
@@ 2.1 Data source and eligibility criteria
For illustrating our joint model proposal, we used a Flatiron Health electronic health record (EHR)-derived de-identified database of patients who were diagnosed with MM in the US between 1st January 2015 and 28th February 2022 (Birnbaum et al., 2020; Ma et al., 2023). During the study period, the de-identified data came from over 280 US cancer clinics (approximately 800 sites of care). The majority of patients in the database originate from community oncology settings; relative community/academic proportions may vary depending on study cohort. Other eligibility criteria are: (i) at least 18 years of age at MM diagnosis, (ii) no longer than 60 days between initial diagnosis and first activity (visit or line of therapy initiation), (iii) more than three months on-treatment before the end of study follow-up, (iv) no malignancies before MM diagnosis, and (v) patients treated with RVd (a three-drug therapy composed of Lenalidomide, Bortezomib, and Dexamethasone) in the first line of therapy, which was oncologist-defined and rule-based, and is one of the most common regimens for patients with newly diagnosed MM (Punke et al., 2017). The last criterion ensures the homogeneity of the selected subpopulation with respect to the initial treatment received. Thus, 1579 patients met all eligibility criteria.
@@ 2.3 Treatment of continuous baseline covariates
Most continuous covariates contain missing data (10%-56%). In order to address this issue, we apply a log transformation to reduce asymmetry, a standardization (z-score) so that their scales are similar, and a mean imputation to replace missing values. Supplementary Appendix A shows a descriptive summary of these covariates before and after being treated. It is worth mentioning that we also performed the analyses without imputation and, although the sample size was greatly reduced, the conclusions were consistent with those using imputation (see Section 5).
@@ 5 MM data analysis
We evaluate the goodness-of-fit of the joint and competing models, considering metrics compatible with the nature of each outcome. For bi-exponential submodels, individual weighted residuals (IWRES) across time by biomarker did not suggest any model misspecification (Desmée et al., 2017), as the points are evenly scattered around the horizontal zero-line (see Supplementary Appendix C). For the categorical submodel and the competing model, class-weighted classification metrics were applied, such as accuracy, precision, recall, and F1-score (see formulas in Supplementary Appendix B). Table 1 shows the percentage of each metric for both models and using a random classifier.
@@
  • Mathematical equivalence in this particular setting (see Supplementary Appendix B).
@@
We observed a clear higher importance for the top three, growth M-spike, baseline FLC, and decay FLC, where all of them presented VI scores that are considerably higher than the others, with minimal overlap between their margins of variation. This means that, in terms of predictive robustness, the growth rate of the M-spike trajectory, the initial FLC value and its decay rate over time are the characteristics that most impact the predicted treatment for the second line of therapy in a subpopulation initially treated with RVd. Growth FLC, decay M-spike, and age occupy positions from 4 to 6 in the ranking, where when analyzing their margins of variation there is no obvious order between them, but their average scores are higher than the subsequent ranked variables. The average score of the subsequent variables is capable of ranking them, but their margins of variation present non-negligible intersections. Supplementary Appendix D shows the VI ranking using the categorical model (no longitudinal information). Except for the shared latent characteristics, the competing model VI ranking produces results very similar to Figure 1.
@@
In Table 2, we highlight ECOG 2+, age and RVd duration, which are significant and present relative risks in opposite directions for Carfilzomib and Pomalidomide compared to the reference treatment. For example, for RVd duration, this means that the relative risk ratio for a one-year increase in RVd treatment time in the first line of therapy is 1.32 (1/0.76) for being treated with Other over Carfilzomib, while this ratio is 1.19 in favor of Pomalidomide when contrasted with Other. Conversely, the significant M-spike and FLC characteristics are consistent for both comparisons. Indeed, increasing the M-spike decay rate favors the choice of Other over Carfilzomib or Pomalidomide. On the other hand, starting RVd treatment with a high FLC value or increasing the growth FLC rate throughout the first line of therapy increases the likelihood of choosing Carfilzomib or Pomalidomide over Other for the second line of therapy. Supplementary Appendix D shows that the inferential results using the categorical model (no longitudinal information) are consistent with those in Table 2, except for the M-spike and FLC characteristics that are not present in the competing model.
@@
Estimated parameters for the bi-exponential M-spike and FLC submodels are provided in Supplementary Appendix C. Population parameters show similar shapes for the average trajectories of M-spike and FLC biomarkers, where M-spike presents an estimated initial value lower than that of FLC, but its growth and decay rates are higher. Residual error and random effects variances indicate greater variability for FLC trajectories. In addition, significant random effects covariances are consistent for both biomarkers, where the random effects for baseline and growth rate are negatively correlated, while baseline and decay rate are positively correlated.
@@
To illustrate individual-specific interpretations using our joint modeling, we randomly selected two patients, named A and B, whose baseline covariates and categorical outcomes are presented in Supplementary Appendix C. Figure 2(a) shows the M-spike and FLC dynamics of patient A, the fit of these trajectories using bi-exponential submodels, and the predicted probability of each treatment option. Initial M-spike was estimated to be close to 25 g/L with a slight decay over the first year followed by an increase until reaching 30 g/L, while FLC presented values below 20 g/L that were well estimated, but with a beginning of trajectory with great uncertainty due to the absence of FLC measurements for this patient over 40 days after starting RVd treatment. Thereby, the predicted probabilities from the categorical submodel suggested Pomalidomide (highest posterior mean) as a treatment for the second line of therapy, which coincides with the physician’s chosen treatment.
@@ 6 Discussion
In conclusion, our study has filled a gap in the modeling literature, shed light on topics related to the clinical practice of MM patients, and connected a popular machine learning strategy to a joint model. In future work, we envision several extensions of the proposed model. For example, in a high-dimensional longitudinal biomarker framework, the robustness to identify significant association parameters is reduced due to potentially strong correlations among biomarkers, but this problem can be overcome using regularization techniques (Andrinopoulou and Rizopoulos, 2016; Sun and Basu, 2024). We chose a VI strategy as a method for ranking the variables, but the Bayesian framework also provides us with spike-and-slab priors as an alternative approach (Islam et al., 2024). In the presence of heterogeneous subgroups, we recommend adapting our proposal to joint latent class models (Proust-Lima et al., 2014; Andrinopoulou et al., 2020). In some scenarios, researchers may be interested in studying the causal effect of treatment choice, which requires a causal inference approach (Rizopoulos et al., 2024). Furthermore, the probability for each treatment can be dynamically updated as new longitudinal patient information becomes available (Rizopoulos, 2011; Andrinopoulou et al., 2017). Finally, machine learning algorithms can be proposed for each submodel, such as ensembling techniques for longitudinal data (Hu and Szymczak, 2023) and boosting methods for multiclass classification (Tanha et al., 2020).
@@ Acknowledgements
D.A. and J.K.B. were supported by the U.K. Medical Research Council grant MC_UU_00002/5 and the collaboration grant jointly funded by Roche and the University of Cambridge. The authors thank the following colleagues for helpful discussions during the research and review stage of this work: Vallari Shah, Mellissa Williamson, Sarwar Mozumder, Madlaina Breuleux, Pascal Chanu, and Chris Harbron. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A. Descriptive summaries for MM data
Appendix B. Goodness-of-fit criteria
@@
Appendix C. Joint model applied to MM data
@@
Appendix D. Competing model applied to MM data

### v3

@@ 1 Introduction
Many authors have extended the standard joint model. For example, some studies modeled multiple longitudinal outcomes and/or competing risks (Hickey et al., 2016; Rué et al., 2017; Rustand et al., 2024), while others incorporated the dependence of spatial random effects into the joint modeling (Martins et al., 2016, 2017; Rappl et al., 2023). There were also proposals that considered categorical longitudinal and survival outcomes (Choi et al., 2015; Chi et al., 2025), two longitudinal submodels, one continuous and one ordinal (Delporte et al., 2025), and joint models of longitudinal continuous and binary data (Wang et al., 2000; Horrocks and van Den Heuvel, 2009; Lu et al., 2016; Zhou et al., 2023). For the latter, to the best of our knowledge, there are currently no extensions available that accommodate more than two categories. This extension introduces at least two additional complexities: (i) more parameters to be estimated, which requires more data, and (ii) imbalanced categories, leading to poor predictive performance (Kolo, 2010).
@@
The adequacy of the bi-exponential submodels is verified through an analysis of individual weighted residuals (Desmée et al., 2017), while the performance of the categorical submodel is evaluated using metrics for multiclass classification (Grandini et al., 2020). Moreover, as a complement to posterior inference, we present a model-agnostic variable importance approach with the aim of improving the transparency, trustability, and interpretability of the results (Casalicchio et al., 2019).
Although our motivation is multiple myeloma disease, our proposal can be extrapolated to other applications where a categorical outcome is explained by baseline and longitudinal variables. Another potential categorical outcome is the hospital to which a patient may be transferred. Note that in both cases, the categorical outcome may be associated with the severity of the disease in each patient, and prognostic factors may assist in choosing the treatment/hospital. In clinical practice, the predictive probability for each categorical outcome can aid physicians in decision-making, especially those who are early in their careers and have limited practical experience.
@@ 2.1 Data source and eligibility criteria
For illustrating our joint model proposal, we used a Flatiron Health electronic health record (EHR)-derived de-identified database of patients who were diagnosed with MM in the US between 1st January 2015 and 28th February 2022 (Birnbaum et al., 2020; Ma et al., 2023). During the study period, the de-identified data came from over 280 US cancer clinics (approximately 800 sites of care). The majority of patients in the database originate from community oncology settings; relative community/academic proportions may vary depending on study cohort. Other eligibility criteria are: (i) at least 18 years of age at MM diagnosis (minimum age for accessing data), (ii) no longer than 60 days between initial diagnosis and first activity (visit or line of therapy initiation) in order to reduce misclassification bias in treatment exposure, (iii) more than three months on-treatment before the end of the study to allow for follow-up time accrual, (iv) no malignancies before MM diagnosis to prevent selected patients from having more aggressive clinical features due to the presence of other types of malignancies, and (v) patients treated with RVd (a three-drug therapy composed of Lenalidomide, Bortezomib, and Dexamethasone) in the first line of therapy, which was oncologist-defined and rule-based, and is one of the most common regimens for patients with newly diagnosed MM (Punke et al., 2017). The last criterion ensures the homogeneity of the selected subpopulation with respect to the initial treatment received. Thus, 1579 patients met all eligibility criteria.
@@ 2.3 Treatment of continuous baseline covariates
Most continuous covariates contain missing data (10%-56%). In order to address this issue, we apply a log transformation to reduce asymmetry, a standardization (z-score) so that their scales are similar, and a mean imputation to replace missing values. Supplementary Appendix A shows a descriptive summary of these covariates before and after being treated. It is worth mentioning that we also performed the analyses without imputation (see Supplementary Appendix B) and, although the sample size was greatly reduced, the conclusions were consistent with those using imputation (see Section 5). However, we acknowledge that our approach to handling missing data is simplistic. It can only be considered suitable under the missing completely at random (MCAR) assumption, and even then, it ignores the uncertainty stemming from the missingness of the data. In more complex scenarios, including missing at random (MAR) and missing not at random (MNAR), we recommend more sophisticated approaches, such as multiple imputation and machine learning techniques (van Buuren, 2021; Emmanuel et al., 2021).
@@ 5 MM data analysis
We evaluate the goodness-of-fit of the joint and competing models, considering metrics compatible with the nature of each outcome. For bi-exponential submodels, individual weighted residuals (IWRES) across time by biomarker did not suggest any model misspecification (Desmée et al., 2017), as the points are evenly scattered around the horizontal zero-line (see Supplementary Appendix D). For the categorical submodel and the competing model, class-weighted classification metrics were applied, such as accuracy, precision, recall, and F1-score (see formulas in Supplementary Appendix C). Table 1 shows the percentage of each metric for both models and using a random classifier.
@@
  • Mathematical equivalence in this particular setting (see Supplementary Appendix C).
@@
We observed a clear higher importance for the top three, growth M-spike, baseline FLC, and decay FLC, where all of them presented VI scores that are considerably higher than the others, with minimal overlap between their margins of variation. This means that, in terms of predictive robustness, the growth rate of the M-spike trajectory, the initial FLC value and its decay rate over time are the characteristics that most impact the predicted treatment for the second line of therapy in a subpopulation initially treated with RVd. Growth FLC, decay M-spike, and age occupy positions from 4 to 6 in the ranking, where when analyzing their margins of variation there is no obvious order between them, but their average scores are higher than the subsequent ranked variables. The average score of the subsequent variables is capable of ranking them, but their margins of variation present non-negligible intersections. Supplementary Appendix E shows the VI ranking using the categorical model (no longitudinal information). Except for the shared latent characteristics, the competing model VI ranking produces results very similar to Figure 1.
@@
In Table 2, we highlight ECOG 2+, age and RVd duration, which are significant and present relative risks in opposite directions for Carfilzomib and Pomalidomide compared to the reference treatment. For example, for RVd duration, this means that the relative risk ratio for a one-year increase in RVd treatment time in the first line of therapy is 1.32 (1/0.76) for being treated with Other over Carfilzomib, while this ratio is 1.19 in favor of Pomalidomide when contrasted with Other. Conversely, the significant M-spike and FLC characteristics are consistent for both comparisons. Indeed, increasing the M-spike decay rate favors the choice of Other over Carfilzomib or Pomalidomide. On the other hand, starting RVd treatment with a high FLC value or increasing the growth FLC rate throughout the first line of therapy increases the likelihood of choosing Carfilzomib or Pomalidomide over Other for the second line of therapy. Supplementary Appendix E shows that the inferential results using the categorical model (no longitudinal information) are consistent with those in Table 2, except for the M-spike and FLC characteristics that are not present in the competing model.
@@
Estimated parameters for the bi-exponential M-spike and FLC submodels are provided in Supplementary Appendix D. Population parameters show similar shapes for the average trajectories of M-spike and FLC biomarkers, where M-spike presents an estimated initial value lower than that of FLC, but its growth and decay rates are higher. Residual error and random effects variances indicate greater variability for FLC trajectories. In addition, significant random effects covariances are consistent for both biomarkers, where the random effects for baseline and growth rate are negatively correlated, while baseline and decay rate are positively correlated.
@@
To illustrate individual-specific interpretations using our joint modeling, we randomly selected two patients, named A and B, whose baseline covariates and categorical outcomes are presented in Supplementary Appendix D. Figure 2(a) shows the M-spike and FLC dynamics of patient A, the fit of these trajectories using bi-exponential submodels, and the predicted probability of each treatment option. Initial M-spike was estimated to be close to 25 g/L with a slight decay over the first year followed by an increase until reaching 30 g/L, while FLC presented values below 20 g/L that were well estimated, but with a beginning of trajectory with great uncertainty due to the absence of FLC measurements for this patient over 40 days after starting RVd treatment. Thereby, the predicted probabilities from the categorical submodel suggested Pomalidomide (highest posterior mean) as a treatment for the second line of therapy, which coincides with the physician’s chosen treatment.
@@ 6 Discussion
Our study has filled a gap in the modeling literature, shed light on topics related to the clinical practice of MM patients, and connected a popular machine learning strategy to a joint model. In future work, we envision several extensions of the proposed model. For example, in a high-dimensional longitudinal biomarker framework, the robustness to identify significant association parameters is reduced due to potentially strong correlations among biomarkers, but this problem can be overcome using regularization techniques (Andrinopoulou and Rizopoulos, 2016; Sun and Basu, 2024). We chose a VI strategy as a method for ranking the variables, but the Bayesian framework also provides us with spike-and-slab priors as an alternative approach (Islam et al., 2024). Standard permutation-based VI methods do not consider the correlation between (baseline and/or longitudinal) variables, which could potentially lead to a misleading ranking; therefore, approaches that take correlation into account should be employed (Nicodemus et al., 2010; Chamma et al., 2024). In the presence of heterogeneous subgroups, we recommend adapting our proposal to joint latent class models (Proust-Lima et al., 2014; Andrinopoulou et al., 2020; Chen et al., 2025). In some scenarios, researchers may be interested in studying the causal effect of treatment choice, which requires a causal inference approach (Rizopoulos et al., 2024). Furthermore, the probability of each treatment can be dynamically updated as new longitudinal patient information becomes available (Rizopoulos, 2011; Andrinopoulou et al., 2017).
In terms of modeling, some alternatives can be considered. For example, the longitudinal submodel can be specified as a GLMM (Brilleman et al., 2018; Alvares and Rubio, 2021), splines (Andrinopoulou et al., 2014), or functional principal components analysis (Guo et al., 2024). Our categorical submodel does not have an explicit time dependence (see Equation (2)), so we cannot share temporal terms from the longitudinal submodel. However, categorical data can be restructured in such a way that at each time point a treatment indicator is observed (see Zhou et al., 2023, for an example with a binary outcome). Specifying the categorical submodel with temporal dependence increases the computational complexity but allows for different association structures, such as the current value of the longitudinal trajectory, its derivative, and its integral (Brown, 2009; Rizopoulos and Ghosh, 2011). Finally, machine learning algorithms can be proposed for each submodel, such as ensembling techniques for longitudinal data (Hu and Szymczak, 2023) and boosting methods for multiclass classification (Tanha et al., 2020).
@@ Acknowledgements
D.A. and J.K.B. were supported by the U.K. Medical Research Council (MRC) grant MC_UU_00002/5, MRC Unit Theme number MC_UU_00040/02 (Precision Medicine), and the collaboration grant jointly funded by Roche and the University of Cambridge. The authors thank the following colleagues for helpful discussions during the research and review stage of this work: Vallari Shah, Mellissa Williamson, Sarwar Mozumder, Madlaina Breuleux, Pascal Chanu, and Chris Harbron. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission.
@@ Appendix A. Descriptive summaries for MM data
Appendix B. Joint model applied to MM data without imputation
Interpretation Parameter M-spike FLC
Baseline $\exp\{\theta_{1kl}\}$ 16.41 (14.35, 20.92) 18.03 (13.67, 25.11)
Growth $\exp\{\theta_{2kl}\}$ 0.17 (0.09, 0.32) 0.15 (0.08, 0.25)
Decay $\exp\{\theta_{3kl}\}$ 5.33 (4.07, 6.17) 4.30 (3.18, 5.25)
Residual error variance $\sigma^{2}_{kl}$ 0.08 (0.04, 0.11) 0.17 (0.12, 0.21)
$\omega_{11kl}$ 0.92 (0.79, 1.02) 3.17 (2.54, 3.64)
Covariance $\omega_{12kl}$0.17 (-0.32, -0.11) -0.98 (-1.62, -0.60)
matrix for $\omega_{13kl}$ 0.22 (0.09, 0.38) 1.22 (0.93, 1.45)
random effects $\omega_{22kl}$ 0.90 (0.73, 1.05) 1.84 (1.57, 2.01)
$\omega_{23kl}$ 0.10 (-0.11, 0.28) -0.07 (-0.26, 0.17)
$\omega_{33kl}$ 0.93 (0.79, 1.09) 2.76 (2.29, 2.98)
Table 6: Relative risk (RR) and its 95% credible interval (CI) from the categorical submodel parameters for Carfilzomib vs Other and Pomalidomide vs Other. Statistically significant variables are shown in bold.
Variable Category Carfilzomib vs Other Pomalidomide vs Other
RR (95% CI) RR (95% CI)
Sex Female 0.92 (0.59, 1.31) 1.00 (0.58, 1.46)
Ethnicity Non-Hisp. Black 1.06 (0.68, 1.52) 0.83 (0.52, 1.37)
Other 1.08 (0.71, 1.83) 0.98 (0.66, 1.69)
ECOG 1 0.65 (0.31, 1.21) 1.12 (0.53, 1.89)
2+ 0.70 (0.43, 0.99) 1.61 (0.99, 2.83)
ISS Stage II 1.56 (0.81, 2.24) 0.62 (0.35, 1.37)
Stage III 1.22 (0.49, 2.08) 1.09 (0.55, 2.44)
Age 0.82 (0.66, 0.94) 1.44 (1.09, 1.62)
Albumin 1.01 (0.83, 1.47) 0.69 (0.55, 0.97)
B2M 1.24 (0.76, 1.61) 0.87 (0.57, 1.40)
Creatine 1.03 (0.75, 1.48) 0.87 (0.59, 1.33)
Hemoglobin 1.01 (0.63, 1.33) 1.09 (0.60, 1.47)
LDH 1.55 (1.14, 1.79) 0.99 (0.60, 1.47)
Lymphocyte 0.88 (0.65, 1.34) 0.79 (0.57, 1.14)
Neutrophil 0.99 (0.77, 1.37) 1.06 (0.71, 1.38)
Platelet 0.98 (0.71, 1.28) 1.04 (0.69, 1.30)
IgA 1.06 (0.74, 1.38) 0.91 (0.69, 1.32)
IgG 1.23 (0.87, 1.55) 0.91 (0.66, 1.48)
IgM 0.89 (0.69, 1.25) 0.79 (0.53, 1.12)
RVd duration 0.61 (0.43, 0.93) 1.15 (0.98, 1.49)
Baseline M-spike 1.20 (0.68, 1.57) 1.18 (0.77, 1.68)
Growth M-spike 1.02 (0.71, 1.83) 1.36 (0.95, 2.58)
Decay M-spike 0.48 (0.25, 0.98) 0.52 (0.21, 0.99)
Baseline FLC 1.55 (1.03, 2.01) 1.73 (1.09, 2.38)
Growth FLC 1.49 (1.08, 2.66) 1.42 (1.01, 2.59)
Decay FLC 0.89 (0.52, 1.17) 0.54 (0.31, 0.93)
Appendix C. Goodness-of-fit criteria
@@
Appendix D. Joint model applied to MM data
@@
Appendix E. Competing model applied to MM data